A 2660 kg satellite used in a cellular telephone network is in a circular orbit at a height of 720 km above the surface of the earth. What is the gravitational force on the satellite?(Answer in N). What fraction is this force of the satellite's weight at the surface on the Earth?(Answer in %).
Answer #1

here

F = G * (m1*m2)/(d^2)

G= gravitational constant (6.67*10^-11 Nm^2/kg^2)
m1 and m2 = the masses object 1 and object 2 respectively
d=the distance between objects

mass of Earth = 5.97 * 10^24 kilograms
mass of satellite = 2660kg
radius of Earth = 6.38 * 10^6 meters
d= 6.38*10^6 meters + 7.20*10^5 meters=7.1 * 10^6 meters

F = 6.67 * 10^-11 * 5.97 * 10^24 * 2660 / ( 7.1 * 10^6)^2

F = 21012 N

then

Ffraction = 21012 / ( 2660 * 9.8 )

Ffraction = 0.8

in percentage it is 80%